OpenStudy (anonymous):

can you help me to find the eqaution,center and radius? : center at origin, circle passes through (4,1)

OpenStudy (mikurout):

general equation is x^2 +y^2 = r^2 for center at origin.

OpenStudy (mikurout):

here radius r is the distance between (0,0) and (4,1).

OpenStudy (mokeira):

First step is to find the radius. You have to points: (0,0) and (4,1) \[\sqrt{[(4-0)^{2}+(1-0)^{2}}\] =\[\sqrt{17}\] so radius= square root of 17 Next, use the fomular\[(x-a)^{2}+(y-b)^{2}=r ^{2}\] \[(x-4)^{2}+(y-1)^{2}=17\] Open up the brackets and...voila! \[x ^{2}-8x+y ^{2}-2y=0\]

OpenStudy (mokeira):

To find the centre: \[[(x _{1}+x _{2})]\div2, [(y _{1}+y _{2})]\div2\] \[([0+4]\div2), ([0+1]\div2)\] (2,0.5) = centre

OpenStudy (anonymous):

Thank you for helping me guys!

OpenStudy (mikurout):

@Mokeira This is not right for the above question. here it is given that center at origin. So the equation of the circle will be x^2 + y^2 = 17

OpenStudy (anonymous):

so the center is (0,0) and raduis is 17? @mikurout

OpenStudy (mikurout):

@SeniorHigh yes, center is ( 0,0) but radius is square root of 17.

OpenStudy (mokeira):

@mikurout Oh yeah! I am sorry! my bad...thats correct

OpenStudy (mokeira):

centre is (0,0) and radius is \[\sqrt{17}\]

OpenStudy (anonymous):

how do you geet the 17?

OpenStudy (mokeira):

@SeniorHigh check in my first answer the explanation for radius is correct

OpenStudy (mikurout):

Use the distance formula and find the distance between the center (0,0) and a point on the circle (4,1).

OpenStudy (mikurout):